OpenStudy (anonymous):

solve the equation on the interval 0<θ <2pie sin(3θ )=-1

OpenStudy (jdoe0001):

\(\bf sin(3\theta) = -1 \implies sin^{-1}\left(sin(3\theta)\right) = sin^{-1}(-1 )\) what would that give you?